OpenStudy (anonymous):

help me with chain rule.... i have u = f(x ) ; x = yt^(-1/2) find d^2u/dy^2.

OpenStudy (amistre64):

looks like we got do some implicit stuff; or... simply regard t as independant; they come out different depending on what we are going for

OpenStudy (anonymous):

can u show e the working... this is using the tree diagram

OpenStudy (anonymous):

i found partial u partial y = du/dx t^(-1/2)

OpenStudy (anonymous):

then from what i get how to find partial^2 u/ partial y^2?

OpenStudy (amistre64):

im not sure what is meant by 'tree' diagram; it sounds rather 'author' specific and not a general mathing term

OpenStudy (anonymous):

what is with the "t"?

OpenStudy (anonymous):

if there is a t in it, but you are taking the derivative of u wrt y isn't the 't' irrelevant?

OpenStudy (amistre64):

depends on how y interacts with y :) right?

OpenStudy (anonymous):

i am just asking, not saying. i am confused.

OpenStudy (amistre64):

\[y = x\sqrt{t}\] right?

OpenStudy (anonymous):

i read it as \[u = f(\sqrt{t}y)\]

OpenStudy (anonymous):

but you want the second derivative wrt y, not t

OpenStudy (anonymous):

no its u = f (y / t^(1/2) )

OpenStudy (anonymous):

yes you are right i didn't see the - sign sorry

OpenStudy (anonymous):

@satelite : yes...scnd derivative wrt to y

OpenStudy (amistre64):

it feels like we are trying to come into this halfway thru the problem; is there a bigger picture that this question fits with?

OpenStudy (anonymous):

still as far as y taking the derivative wrt y \[\frac{1}{\sqrt{t}}\] is a constant

OpenStudy (anonymous):

the tree diagram that i was referring to is like this http://www.youtube.com/watch?v=aZcw1kN6B8Y

OpenStudy (anonymous):

and we don't know what u is. \[u=f(x)\] tells us nothing. so it is just \[u'=f'(\frac{y}{\sqrt{t}})\frac{1}{\sqrt{t}}\]

OpenStudy (anonymous):

ok got that already

OpenStudy (anonymous):

is there more to the problem than written? because this looks nothing like taking partials as in the video

OpenStudy (anonymous):

im sorry i dont understand... i am looking for a partials 2nd derivative...from the video is only 1st derivative.. so how do i continue from the answer u gave to get second partial wrt to y